OpenStudy (anonymous):

I NEED HELP PLZ HELP.........You are making scarves for each presents. Each scarf needs 3/4 yd of fabric. How many yards of fabric do you need for 6 scarves? A. 9 1/3 B. 4 1/2 C. 8 yd D.5 1/4

OpenStudy (anonymous):

I think its B. 4 1/2

OpenStudy (retireed):

Show me some ciphering.

OpenStudy (anonymous):

I dont know what you mean

RhondaSommer (rhondasommer):

what is 6 * 3/4?

OpenStudy (retireed):

Google cipher or watch to Beverly Hillbillies about Jethro and is powerful math skills.

RhondaSommer (rhondasommer):

6 3 - * - 1 4

RhondaSommer (rhondasommer):

multiply across :)

RhondaSommer (rhondasommer):

6*3 = ?

OpenStudy (anonymous):

6*3/4 is 4 1/2

RhondaSommer (rhondasommer):

you are correct! :)

OpenStudy (anonymous):

:)

RhondaSommer (rhondasommer):

glad i could help. any more questions?

OpenStudy (anonymous):

no thank you so much

OpenStudy (tkhunny):

Make your life simpler... \(6\cdot \dfrac{3}{4} = \dfrac{6}{4}\cdot 3 = \dfrac{3}{2}\cdot 3\)